# LBL: Logarithmic Barrier Loss Function for One-class Classification

### v2

@@ Abstract
One-class classification (OCC) aims to train a classifier solely on target data and attracts increasing attention due to its applicability in practice. Despite OCC has obtained many advances, it still lacks the effective OCC loss functions for deep learning. In this paper, a novel logarithmic barrier function based OCC loss (LBL) that assigns large gradients to margin samples and thus derives more compact hypersphere is first proposed by approximating the OCC objective smoothly. But the optimization of LBL may be instability especially when samples lie on the boundary leading to the infinity value. To address this issue, a smoother LBLSig loss is further proposed by utilizing a unilateral relaxation Sigmoid function. Experiments on different networks demonstrate the effectiveness of the proposed LBL and LBLSig. The source code can be found at https://github.com/ML-HDU/LBL_LBLSig.
Index Terms—  One-class classification, anomaly detection, logarithmic barrier loss, hypersphere optimization
1 Introduction
One-class classification (OCC), also known as the anomaly detection (AD), aims to construct the model only exploiting the characteristics of target data. It has attracted much attention in various communities such as data mining, machine learning and computer vision [1]. The classical methods [2, 3] are generally required feature engineering. In comparisons, deep learning-based OCC algorithms achieve promising performance including the autoencoder (AE) based methods [4, 5, 6, 7, 8], the generative adversarial network (GAN) based methods [9, 10, 11], and other discrimination-based methods [12, 13, 14, 15, 16].
However, little attention is paid to the OCC loss function design. Recently typical OCC loss functions are developed in [13] and [14]. Ruff et al. [13] proposes the soft-boundary loss (SBL) that can be regarded as an extension of the mean square error (MSE) in OCC. But SBL is only demonstrated on the LeNet-type convolution neural network (CNN) and the performance is not promising. In addition, SBL is essentially the hard sample mining with hinge function form that is non-smooth and difficult to optimize. Different from the MSE-type SBL, a cross entropy (CE)-type loss function named HRN that consists of a negative log-likelihood (NLL) term and a gradient penalty term with respect to inputs is proposed in [14]. The HRN shows the surprising results but the performance both in theory and experiment is only verified on the simple multilayer perceptron (MLP). Particularly, we find that the HRN is optimized difficultly in practice and the performance is poor on CNN. Therefore, it still lacks effective OCC loss functions for deep learning.
@@ 1 Introduction
Actually, OCC aims to seek a hypersphere with the minimum radius to enclose the target data. In this paper, a novel logarithmic barrier function [17] based OCC loss function (LBL) is first proposed by approximating the OCC objective smoothly. LBL assigns larger gradients to these samples close to boundary, and thus achieves more compact hypersphere. However, the direct approximation by the logarithmic barrier function may cause the instability especially when samples lie on the boundary leading to the infinity loss. To address this issue, then, a unilateral relaxation Sigmoid function is introduced into LBL and a novel OCC loss named LBLSig is further proposed. Owing to the unilateral relaxation Sigmoid function, the optimization becomes smoother. Particularly, LBLSig can be seen as the fusion of MSE and CE. Experiments on various networks are conducted to verify the effectiveness of the proposed LBL/LBLSig. The results demonstrate the superior performance of our proposed algorithms.
2 Problem Statement
@@
where $\lambda_{1}$ and $\lambda_{2}$ are the trade-off parameters, $R$ is the radius of hypersphere. It is worth pointing out that the radius is updated in each mini-batch [6] by choosing a certain quantile from the sequence $\left\|{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)-{\bf{c}}}\right\|^{2}$, as shown in the implementation of SBL [^1] Code of Deep SVDD: https://github.com/lukasruff/Deep-SVDD-PyTorch.. Therefore, SBL is the iterative hard sample mining in essence.
@@
$\small\min_{\boldsymbol{\mathcal{W}}}\quad\underbrace{-\sum_{i=1}^{N}{{\log\left({{\text{Sig}}\left({{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}\right)}\right)}}}_{\text{NLL}}+{\lambda}\underbrace{\sum_{i=1}^{N}{\left\|{{\nabla}_{\bf{x}}{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}}\right\|_{F}^{q}}}_{\text{H-regularization}},$ (3)
where $\lambda$ and $q$ are two manual hyperparameters, and $\text{Sig}(\cdot)$ is the Sigmoid function that can be seen as the probability belonging to targets. It can be readily seen that minimizing NLL means to obtain high output value for target data, and the NLL term is essentially the direct extension from CE. H-regularization term is introduced as a constraint over the growth on the value of ${\boldsymbol{\Phi}}\left({\bf{x}}\right)$} caused by the single target class. HRN shows the superior performance than SBL but it is only verified on MLP. In practice, we find that the HRN is optimized difficultly and the performance is poor on CNN.
3 Proposed methods
3.1 LBL
@@
Obviously, minimizing $I(u)$ can enforce ${\boldsymbol{\Phi}}({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}})$ to fall inside the hypersphere $({\bf{c}},{R})$. However, (5) is difficult to be optimized due to that $I(u)$ is not differentiable. To address this issue, LBL is proposed to approximate (6) smoothly by the logarithmic barrier function [17] as
@@ 3.2 LBLSig
${\frac{\partial{J_{2}}}{\partial{\boldsymbol{\mathcal{W}}}}}$ $={\frac{1}{N}}\sum_{i=1}^{N}{\left({1-v_{i}}\right){\frac{\partial{D^{2}_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}.$ (12)
@@
4 Experiments
4.1 Implementation details
The proposed LBL and LBLSig are conducted on $4$ different networks for comparisons with SBL [13] and HRN [14], including DSLeNet [13], MLP with $2$ hidden layers, and OCITN [15]. For DSLeNet, the center $\bf{c}$ is set to be the mean of the outputs of the initialized DSLeNet on training data. For MLP, the mean of the training samples is computed as the hypersphere center $\bf{c}$. For OCITN, the “Lena image” is chosen as the center $\bf{c}$ the same as [15] (HRN is hence inapplicable to OCITN). The grid search method is employed to select the optimal hyperparameters for all algorithms. More detailed settings can be found in source code due to the page limitation. In addition, SBL argues in [13] that networks with bias will lead to hypersphere collapse. Actually, SBL is non-convex and the network with bias leading to hypersphere collapse is a small probability. We hence do not add any constrains to the bias term in this paper.
Table 1: Comparisons of AUCs with SBL and HRN on CIFAR10.
@@ 4.1 Implementation details
Table 3: Comparisons of AUCs with the SOTA OCC algorithms.
Target LSAR OCGAN HLS-OC MOCCA HRN LBL LBLSig
Class [7] [10] [8] [6] [14]
AIRPLANE 73.5 75.7 73.7 66.0 77.3 76.38 78.03
AUTOMOBILE 58.0 53.1 74.4 74.6 69.9 87.93 87.29
BIRD 69.0 64.0 60.9 57.5 60.6 68.82 69.42
CAT 54.2 62.0 63.9 60.1 64.4 60.50 60.91
DEER 76.1 72.3 71.0 61.5 71.5 71.01 71.18
DOG 54.6 62.0 64.1 68.4 67.4 72.17 71.99
FROG 75.1 72.3 78.8 67.4 77.4 79.83 78.54
HORSE 53.5 57.5 70.6 72.1 64.9 80.82 81.59
SHIP 71.7 82.0 81.8 79.2 82.5 84.33 84.80
TRUCK 54.8 55.4 79.0 77.3 77.3 84.36 84.96
AVG 64.05 65.63 71.82 68.61 71.32 76.62 76.87
  • *Conducted on OCITN
4.2 Comparisons with SBL and HRN
Table 1 gives comparisons of AUCs on CIFAR10 where the best results are highlighted in bold font.
DSLeNet as backbone. The results of SBL are taken from [13] for comparisons. The proposed LBLSig obtains the highest AUC value on average. Specially, LBLSig outperforms the SBL [13] on all classes, and generally has superior performance than HRN.
MLP as backbone. LBL and LBLSig obtain the first and second best performance on average, respectively, and much higher performance than HRN in our MLP structure. Particularly, the results in [14] are obtained using much more complicated MLP than ours. Table 3 lists the results in [14] but LBL and LBLSig still obtain better performance than HRN of [14], while only 2-hidden-layer MLP is used by us.
OCITN as backbone. LBLSig performs the best on 7 out of 10 classes and achieves the highest average AUC. It should be pointed out that HRN is hard to be trained on OCITN, and thus the results fails to be given. In other words, LBL and LBLSig have wider applicability than HRN. In addition, comparisons among networks show that LBL/LBLSig obtains the best performance when using OCITN as backbone.
At last, comparisons on non-image datasets are provided in Table 2, where the single hidden layer MLP is used. Obviously, LBL/LBLSig outperform SBL and HRN on all non-image datasets and LBLSig wins the best average AUC value.
4.3 Comparisons with SOTA algorithms
Comparisons on CIFAR10 with common state-of-the-art (SOTA) OCC algorithms, including LSAR [7], OCGAN [10], HLS-OC [8], MOCCA [6] and HRN [14], are reported. OCITN is chosen as the backbone to conduct LBL/LBLSig. Results are listed in Tabel 3 and it can be seen that the best performance is obtained by the proposed LBL/LBLSig. Specially, LBLSig achieves more than $12\%$, $10\%$, $4\%$, $8\%$ and $5\%$ increments of the average AUC over the LSAR, OCGAN, HLS-OC, MOCCA and HRN, respectively.
5 Conclusions
Table 2: Comparisons of AUCs on non-image datasets.

### v3

@@ Abstract
One-class classification (OCC) aims to train a classifier only with the target class data and attracts great attention for its strong applicability in real-world application. Despite a lot of advances have been made in OCC, it still lacks the effective OCC loss functions for deep learning. In this paper, a novel logarithmic barrier function based OCC loss (LBL) that assigns large gradients to the margin samples and thus derives more compact hypersphere, is first proposed by approximating the OCC objective smoothly. But the optimization of LBL may be instability especially when samples lie on the boundary leading to the infinity loss. To address this issue, then, a unilateral relaxation Sigmoid function is introduced into LBL and a novel OCC loss named LBLSig is proposed. The LBLSig can be seen as the fusion of the mean square error (MSE) and the cross entropy (CE) and the optimization of LBLSig is smoother owing to the unilateral relaxation Sigmoid function. The effectiveness of the proposed LBL and LBLSig is experimentally demonstrated in comparisons with several popular OCC algorithms on different network structures. The source code can be found at https://github.com/ML-HDU/LBL_LBLSig.
I Introduction
One-class classification (OCC), also known as the anomaly detection (AD), aims to construct the model only exploiting the characteristics of the target data. It has attracted much attention in various communities, such as data mining, machine learning and computer vision [17, 36], due to its significance in many applications [14, 30]. The classical methods, such as the one-class support vector machine (OCSVM) [27] and the support vector data description (SVDD) [28], are confronted with difficulties when dealing with high-dimensional and complex data, and such shallow methods are always required feature engineering. In contrast, deep learning-based OCC algorithms achieve exceptional successes, including the autoencoder (AE) based methods [19, 37, 11, 33, 24, 3, 6, 16, 1, 32, 4, 34, 9], the generative adversarial network (GAN) based methods [26, 22, 35, 2, 23, 18, 25], and other discrimination-based methods [10, 29, 21, 13, 12, 20, 7]. However, little attention is paid to the OCC loss function design. Recently typical OCC loss functions are developed in [21] and [13]. Ruff et al. [21] propose the soft-boundary loss (SBL) that can be regarded as an extension of the mean square error (MSE) in OCC. But SBL is only demonstrated on the LeNet-type convolution neural network (CNN) and the performance is not promising. In addition, SBL is essentially the hard sample mining with non-smooth hinge function form. Different from the MSE-type SBL, a cross entropy (CE)-type loss function named HRN that consists of a negative log-likelihood (NLL) term and a gradient penalty term with respect to inputs is proposed in [13]. OCC is indeed a binary classification obeying the Bernoulli distribution. However, the biggest difference from the binary classification is that the non-target samples are absent in training. The training of cross entropy solely on target samples may result in malformed learnable parameters. The gradient regularization is employed to alleviated this issue, but the training becomes harder. In addition, HRN is only verified on the simple multilayer perceptron (MLP).
@@ I Introduction
In this paper, we follow the OCC design that seeks a hypersphere with the minimum radius to enclose the target data. Specifically, a novel logarithmic barrier function based OCC loss function (LBL) is first proposed by approximating the OCC objective smoothly. LBL assigns larger gradients to samples close to boundary, and thus achieves more compact hypersphere. However, the direct approximation by the logarithmic barrier function may cause the instability when samples lie on the boundary leading to the infinity loss. To address this issue, then, a unilateral relaxation Sigmoid function is introduced into LBL and a novel OCC loss named LBLSig is further proposed. Owing to the unilateral relaxation Sigmoid function, the optimization becomes smoother. Particularly, LBLSig can be seen as the fusion of MSE and CE. Experiments on various networks are conducted to verify the effectiveness of the proposed LBL/LBLSig. The results demonstrate the superior performance of our proposed algorithms.
II Brief Review
@@
where $\lambda_{1}$ and $\lambda_{2}$ are the trade-off parameters, $R$ is the radius of hypersphere. It is worth pointing out that the radius is updated in each mini-batch [16] by choosing a certain quantile from the sequence $\left\|{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)-{\bf{c}}}\right\|^{2}$, as shown in the implementation [^1] Code of Deep SVDD: https://github.com/lukasruff/Deep-SVDD-PyTorch.. Therefore, SBL is the iterative hard sample mining in essence.
@@
$\min_{\boldsymbol{\mathcal{W}}}\quad\underbrace{-\sum_{i=1}^{N}{{\log\left({{\text{Sig}}\left({{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}\right)}\right)}}}_{\text{NLL}}+{\lambda}\underbrace{\sum_{i=1}^{N}{\left\|{{\nabla}_{\bf{x}}{{\boldsymbol{\Phi}}\left({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}}\right)}}\right\|_{F}^{q}}}_{\text{H-regularization}},$ (3)
where $\lambda$ and $q$ are two manual hyperparameters, and $\text{Sig}(\cdot)$ is the Sigmoid function that can be seen as the probability belonging to targets. It can be readily seen that minimizing NLL means to obtain high output value for target data, and the NLL term is essentially the direct extension from CE. H-regularization term is introduced as a constraint over the growth on the value of ${\boldsymbol{\Phi}}\left({\bf{x}}\right)$} caused by the single target class.
III Proposed methods
III-A LBL
@@
Obviously, minimizing $I(u)$ can enforce ${\boldsymbol{\Phi}}({\bf{x}}_{i};{\boldsymbol{\mathcal{W}}})$ to fall inside the hypersphere $({\bf{c}},{R})$. However, (5) is difficult to be optimized due to that $I(u)$ is not differentiable. To address this issue, LBL is proposed to approximate (6) smoothly by the logarithmic barrier function [5, 38] as
@@ III-B LBLSig
${\frac{\partial{J_{2}}}{\partial{\boldsymbol{\mathcal{W}}}}}$ $=\sum_{i=1}^{N}{{\frac{\partial{J_{2}}}{\partial{v_{i}}}}{\frac{\partial{v_{i}}}{\partial{u_{i}}}}{\frac{\partial{u_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}$ (12)
$={\frac{1}{N}}\sum_{i=1}^{N}{{\frac{e^{u_{i}}}{1+e^{u_{i}}}}{\frac{\partial{D^{2}_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}$
$={\frac{1}{N}}\sum_{i=1}^{N}{\left({1-v_{i}}\right){\frac{\partial{D^{2}_{i}}}{\partial{\boldsymbol{\mathcal{W}}}}}}.$
@@
IV Experiments
TABLE I: Comparisons of AUCs with SBL and HRN on CIFAR10.
@@ IV Experiments
TABLE II: Comparisons of AUCs on non-image datasets.
@@
IV-A Implementation details
The proposed LBL and LBLSig are conducted on $4$ different networks for comparisons with SBL [21] and HRN [13], including DSLeNet [21], MLP with $2$ hidden layers, and OCITN [12]. For DSLeNet, the center $\bf{c}$ is set to be the mean of the outputs of the initialized DSLeNet on training data. For MLP, the mean of the training samples is computed as the hypersphere center $\bf{c}$. For OCITN, the center $\bf{c}$ is determined the same as [12]. The grid search method is employed to select the optimal hyperparameters for all algorithms. More detailed settings can be found in source code due to the page limitation. In addition, SBL argues in [21] that networks with bias will lead to hypersphere collapse. Actually, SBL is non-convex and the network with bias leading to hypersphere collapse is a small probability. We hence do not add any constrains to the bias term in this paper.
TABLE III: Comparisons of AUCs with the SOTA OCC algorithms.
Target LSAR OCGAN HLS-OC MOCCA HRN LBL LBLSig
Class [1] [18] [32] [16] [13]
AIRPLANE 73.5 75.7 73.7 66.0 77.3 76.38 78.03
AUTOMOBILE 58.0 53.1 74.4 74.6 69.9 87.93 87.29
BIRD 69.0 64.0 60.9 57.5 60.6 68.82 69.42
CAT 54.2 62.0 63.9 60.1 64.4 60.50 60.91
DEER 76.1 72.3 71.0 61.5 71.5 71.01 71.18
DOG 54.6 62.0 64.1 68.4 67.4 72.17 71.99
FROG 75.1 72.3 78.8 67.4 77.4 79.83 78.54
HORSE 53.5 57.5 70.6 72.1 64.9 80.82 81.59
SHIP 71.7 82.0 81.8 79.2 82.5 84.33 84.80
TRUCK 54.8 55.4 79.0 77.3 77.3 84.36 84.96
AVG 64.05 65.63 71.82 68.61 71.32 76.62 76.87
  • *Conducted on OCITN
IV-B Comparisons with SBL and HRN
Table I gives the comparisons of AUCs on CIFAR10 under 3 different networks where the best results are highlighted in bold font.
DSLeNet as backbone. The results of SBL are taken from [21] for comparisons. The proposed LBLSig obtains the highest AUC value on average. Specially, LBLSig outperforms the SBL [21] on all classes and the increments are $10.98\%$ on AIRPLANE, $7.05\%$ on HORSE and $4.8\%$ on average. Similarly, the LBLSig generally has superior performance than HRN except the AUTOMOBILE class, and achieves more than $14.41\%$ increments on DOG, $18.72\%$ increments on HORSE and $7.56\%$ increments on average.
MLP as backbone. It can be readily seen that LBL and LBLSig obtain the first and second best performance on average, respectively. Particularly, LBL and LBLSig obtain much higher performance than HRN in our MLP structure. But it is worthy pointing out that the results reported in [13] are obtained using the structure of $\{3\times[1024$-300]}300]\}-[900[900-300]300]-[300[300-$1]$ which is more complicated than our 2-hidden-layer MLP. The results in [13] are listed in Table III, and it can be readily seen by comparisons that LBL and LBLSig generally outperform HRN of [13] and obtain better average AUC while only 2-hidden-layer MLP is used by us.
OCITN as backbone. As shown in Table I, LBLSig performs the best on 7 out of 10 classes and achieves the highest average AUC. On the contrary, the SBL on OCITN fails to obtain the promising performance and LBLSig has average $12.83\%$ improvements over SBL. In addition, HRN is originally required to fit a scalar, but OCITN aims to a color image, which leads to the hard optimization of HRN on OCITN. The results are hence fails to given in the table. However, it can be also said that LBL and LBLSig have wider applicability than HRN. In addition, the comparisons among networks show that LBL/LBLSig obtains the best performance when using OCITN as backbone. Particularly, LBLSig on OCITN achieves $4.49\%$ and $8.79\%$ increments over DSLeNet and MLP, respectively.
It should be pointed out that LBL may suffer from instability optimization and low robustness and thus LBLSig is developed to improve stability and robustness. LBLSig performs a litter better than LBL in experimental comparisons because the experiments are conducted on clean data such as CIFAR10. Due to the page limitation, the comparisons of robustness will demonstrated in future work.
At last, the comparisons on non-image datasets are conducted and shown in Table II, where the single hidden layer MLP is used as the backbone. It can be readily seen that LBL/LBLSig outperform SBL and HRN on all non-image datasets and LBLSig wins the best average AUC value.
IV-C Comparisons with SOTA algorithms
Comparisons on CIFAR10 with common state-of-the-art (SOTA) OCC algorithms, including LSAR [1], OCGAN [18], HLS-OC [32], MOCCA [16] and HRN [13], are reported. OCITN is chosen as the backbone to conduct LBL/LBLSig. Results are listed in Tabel III and it can be seen that the best performance is obtained by the proposed LBL/LBLSig. Specially, LBLSig achieves more than $12\%$, $10\%$, $4\%$, $8\%$ and $5\%$ increments of the average AUC over the LSAR, OCGAN, HLS-OC, MOCCA and HRN, respectively.
V Conclusions